OpenStudy (cutiecomittee123):

which if the following are identities sin3x/sinxcosx=4cosx-secx (sinx+cosx)^2=1+sin2x sin3x-sinx/cos3x+cosx=tanx sin6x=2sin3xcos3x

OpenStudy (tkhunny):

The first is ambiguous and should be written more carefully. Use your best known identities and prove them!

OpenStudy (cutiecomittee123):

Well just a wild guess but the second one looks like one

OpenStudy (tkhunny):

Wild guesses are not acceptable. Prove it. Expand the left side and see.

OpenStudy (cutiecomittee123):

I am not sure how that is done

OpenStudy (tkhunny):

The last is an eyeball question. Just look at it, recall the double angle formula for sine, and call it good.

OpenStudy (tkhunny):

Not sure how what is done? It's a binomial. Square it.

OpenStudy (cutiecomittee123):

I see how the last one works though since the double identity is sin2x=2sinxcosx and it says sin6x=2sin3xcos3x

OpenStudy (cutiecomittee123):

I am still not seeing any other ones

OpenStudy (tkhunny):

Done. Now, #2 (a+b)^2 = a^2 + 2ab + b^2 Apply this idea to the left side.

OpenStudy (cutiecomittee123):

(sinx+cosx)^2=1+sin2x sinx+cosx)^2=sin^2x+2sinxcosx+cos^2x

OpenStudy (cutiecomittee123):

now what though?

OpenStudy (tkhunny):

You're done. Just look at it. sin^2(x) + cos^2(x) = 1 That should look familiar; Just play with them. They won't break.